OpenStudy (anonymous):

The height in feet reached by a baseball after t seconds is given by: H(t)=-16t^2+88t+4. Write and solve an equation to determine when the baseball is 100 ft in the air.

OpenStudy (anonymous):

How would I approach this problem?

OpenStudy (anonymous):

Substitute 100 in for H 100= -16t^2+88t+4

OpenStudy (aum):

Correct. Add +16t^2 - 88t - 4 to both sides. Simplify. Solve the quadratic. You will get two values for t. That means the baseball reaches the height 100 feet once while going up and then again when coming down.

OpenStudy (anonymous):

If I wanted to use a graphing calculator to solve this equation how would I approach it?

OpenStudy (anonymous):

Set Y=-16t^2+88t -96 I think then what else

OpenStudy (aum):

100= -16t^2+88t+4 Add +16t^2 - 88t - 4 to both sides: 100 +16t^2 - 88t - 4 = 0 16t^2 - 88t + 96 = 0 (same as your equation after multiplying by -1) You can set y =16t^2 - 88t + 96. The solution using the graphing calculator will be the time where the graph crosses the x-axis where y will be zero. And we are trying to solve the equation when y will be zero.

OpenStudy (anonymous):

Wouldnt it be -96?

OpenStudy (aum):

-16t^2+88t -96 = 0 multiply by -1: 16t^2 - 88t + 96 = 0

OpenStudy (aum):

Or if you want you can just plot -16t^2 + 88t - 96 You will get the same answer.